# AmpLyze: A Deep Learning Model for Predicting the Hemolytic Concentration

### v1

@@ 2 RESULTS
Hyperparameters have a direct impact on model performance, so careful tuning is essential. We performed a grid search across the hyperparameter ranges listed in Table 3. In experiments, we discovered that the loss function had the greatest influence on the results. Because the loss guides gradient descent and the learning process, it is arguably the most critical hyperparameter in deep-learning models. In our context of hemolytic concentration, it is worth noting that HC50 values are notoriously protocol-sensitive. The same peptide, for example, can produce hemolysis ratios that vary by up to four-fold when the assay is run on mouse versus rabbit erythrocytes, and simply substituting the positive-control detergent can alter the value by a further approximately three-fold [25]. Our dataset comprises HC50 values measured with human, mouse, horse and rabbit erythrocytes, yet key protocol details are often left unspecified; consequently, substantial inter-experiment variability is inevitable. To limit overfitting and ensure generalization across these heterogeneous conditions, a robust loss function is fundamentally essential. We evaluated Mean Square Loss (MSE), Mean Absolute Loss (MAE), Huber Loss and Log-Cosh Loss (Eq. 3). As expected, MSE performed worst because of its pronounced sensitivity to outliers. Huber Loss comes with an additional hyperparameter to tune to control the switch from quadratic and linear loss. We experimented with $\delta=0.5$ and $\delta=1.0$, with both giving better results than MAE, but still slightly trailed Log-Cosh. Log-Cosh seamlessly integrates quadratic and linear penalties into a single smooth, twice-differentiable function, eliminating the need to hand-tune a $\delta$ hyperparameter while dampening the influence of large residuals.
@@
Table 3: Hyperparameter search space for AmpLyze.
                Hyperparameter         Candidate values         Best value
@@
Figure 3: Surface hydrophobicity and backbone representation of wild-type and mutant Temporin peptides. (A) Wild-type Temporin shows a continuous hydrophobicity surface with the flexible C-terminus. (B) Single mutant Temporin-K12 (substituion in magenta) disrupts the hydrophobicity surface by introducing a positively charged side chain. (C) Triple mutant Temporin-4K breaks up nonpolar clustering with three lysine substitutions. Peptide 3D structures were predicted using PEP-FOLD 3.5 with default parameters [39].
2.2 Ablation study
To understand the importance of each component of the architecture (Figure 4), we performed an ablation study. After removing one module at a time, we retrained the network while keeping all other settings unchanged, such as dataset spilts, learning-rate schedule and optimizer, and compared the results with the performance of the original model. The three variants are: (1) Global-only, which removes the local branch. The model bases its prediction exclusively on the per-sequence information. (2) Local-only, which strips the global branch, forcing the model to rely solely on per-residue information and ignoring the coarse sequence-level context information. (3) No Cross-attention, which retains both branches but the global-to-local cross-attention mechanism is replaced with simple mean-pooling over the local features. All ablated models were trained from scratch under identical hyper-parameter settings to ensure fair comparison. As summarized in Table 4, dropping either representation branch degrades performance: Local-only (PCC 0.711, MSE 1.195) and Global-only (0.725, 1.107) each lose around 4-6% PCC relative to the full model, confirming that residue-level and sequence-level cues are complementary. Adding cross-attention on top of the dual branches in place of simple pooling elevates performance to PCC 0.756 and MSE 0.987, providing a further 3% reduction in MSE and 1% gain in PCC achieved by dynamically aligning global context with critical residues.
Table 4: Performance of AmpLyze ablation variants
Model PCC $\uparrow$ MSE $\downarrow$ MAE $\downarrow$ R2 $\uparrow$
AmpLyze without Global Branch 0.725 $\pm$ 0.026 1.107 $\pm$ 0.108 0.744 $\pm$ 0.032 0.518 $\pm$ 0.04
AmpLyze without Local Branch 0.711 $\pm$ 0.023 1.195 $\pm$ 0.101 0.773 $\pm$ 0.028 0.471 $\pm$ 0.045
AmpLyze without Cross Attention 0.748 $\pm$ 0.021 1.014 $\pm$ 0.116 0.716 $\pm$ 0.033 0.550 $\pm$ 0.046
AmpLyze 0.756 $\pm$ 0.019 0.987 $\pm$ 0.095 0.703 $\pm$ 0.029 0.570 $\pm$ 0.033
3 MATERIALS AND METHODS
Data collection.
We employed the same dataset as HemoPI2 to ensure a fair comparison of model performance. All peptide sequences were obtained from two publicly available sources: Hemolytik [40] and DBAASP v3 [41]. Hemolytik is a comprehensive resource that aggregates experimentally validated hemolytic and non-hemolytic peptides from several well-established antimicrobial peptide (AMP) databases, including APD2 and DAMPD, while DBAASP also contains detailed activity profiles that includes hemolytic and cytotoxic activity. Only AMP entries with available and numeric HC50 values were retained; for peptides with multiple HC50 measurements or reported ranges, the mean value was used to represent overall hemolytic activity. Sequences containing non-standard residues or shorter than six amino acids were removed. After filtering, the final dataset comprised 1926 peptides.
To train and evaluation our models, the entire dataset was partitioned into five mutually exclusive using stratified sampling that preserved the peptide-length and HC50 distribution to limit potential distribution shift. In each iteration, four folds were used to train the model and tune hyper-parameters, while the remaining fold served as an unseen validation set to prevent data leakage. Model performance is then reported as the mean $\pm$ standard deviation across the five runs.
$\text{pHC}_{50}=-\ln\bigl{(}\text{HC}_{50}\bigr{)}$ (4)
Data augmentation.
Evaluation metrics.
We employed mean squared error (MSE), mean absolute error (MAE), Pearson correlation coefficient (PCC) and coefficient of determination ($R^{2}$) to measure the performance of our model. They are defined as follows:
$\mathrm{MSE}=\frac{1}{N}\sum_{i=1}^{N}\bigl{(}y_{i}-\hat{y}_{i}\bigr{)}^{2}$ (5)
$\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}\bigl{|}y_{i}-\hat{y}_{i}\bigr{|}$ (6)
$\mathrm{PCC}=\frac{\sum_{i=1}^{N}(y_{i}-\bar{y})(\hat{y}_{i}-\overline{\hat{y}})}{\sqrt{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}\,\sqrt{\sum_{i=1}^{N}(\hat{y}_{i}-\overline{\hat{y}})^{2}}}$ (7)
$R^{2}=1-\frac{\displaystyle\sum_{i=1}^{N}\bigl{(}y_{i}-\hat{y}_{i}\bigr{)}^{2}}{\displaystyle\sum_{i=1}^{N}\bigl{(}y_{i}-\bar{y}\bigr{)}^{2}}$ (8)
where
$\bar{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i},\qquad\overline{\hat{y}}=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}.$
Figure 4: The proposed AmpLyze architecture based on (A) local and global encoding branches and enhanced with (B) a cross-attention fusion module.
Pre-trained protein language models.
To encode rich biochemical context from raw amino acid sequences, we used two state-of-the-art pre-trained protein language models: ESM2-3B [43] and ProtT5-XL-UniRef50 [44]. The ESM2 3B model comprises 36 transformer layers and roughly 3 billion parameters, trained on over 250 million UniRef50 sequences with a masked-language modeling objective. We input the unmodified protein sequence and extracted the final hidden-state embeddings for each residue. ProtT5 is an encoder-only variant of the T5 architecture with 24 layers and 3 billion parameters, trained on UniRef50 in half precision. From ProtT5 we extracted both per-residue embeddings and a per-sequence embedding obtained by mean-pooling across all residues. In our framework, per-residue embeddings were fed into the local branch, while per-sequence embeddings were used in the global branch of AmpLyze.
4 Discussion
@@
In this work, we present AmpLyze, a sequence-based deep learning model to predict peptide hemolytic concentration (HC50) quantitatively and interpretably. By leveraging pre-trained protein language model embeddings and a fusion-embedding strategy, AmpLyze achieves superior performance over prior regression approaches: Pearson’s $r$ improved from 0.739 to 0.756, $R^{2}$ rose from 0.543 to 0.570, MSE fell from 1.074 to 0.987, and MAE declined from 0.753 to 0.703. These gains demonstrate that deep neural architectures can more effectively capture the nonlinear, high-dimensional relationships between sequence motifs and hemolytic activity. We further show that robust loss functions are crucial for handling the inherent heterogeneity of hemolysis assays. Experimental HC50 values can vary by orders of magnitude across different erythrocyte sources, buffer compositions, and protocols. We found that losses such as Log-Cosh enables AmpLyze to generalize across noisy, multi-source datasets. Importantly, AmpLyze offers residue-level interpretability via Expected Gradients with a multi-baseline scheme anchored by non-hemolytic references. This approach successfully recapitulates literature-reported mutational effects, shedding light on rational peptide design.
@@
Looking ahead, the next step is to integrate AmpLyze with in-silico MIC predictors to create a unified HC50–MIC framework. Both hemolytic and antimicrobial activities depend on peptide–membrane interactions: increased cationicity and amphipathicity promote binding to negatively charged bacterial membranes but also enhance insertion into zwitterionic erythrocyte bilayers. Future work will focus on training a multi-task model to capture subtle sequence features that maximize bacterial pore formation while minimizing red-cell lysis. Such a joint optimization strategy promises to accelerate the design of peptides with balanced safety and efficacy.

### v2

@@ III-E Hyperparameter Tuning
Hyperparameters have a direct impact on model performance, so careful tuning is essential. We performed a grid search across the hyperparameter ranges listed in Table IV. In experiments, we discovered that the loss function had the greatest influence on the results. Because the loss guides gradient descent and the learning process, it is arguably the most critical hyperparameter in deep‑learning models. In our context of hemolytic concentration, it is worth noting that HC50 values are notoriously protocol‑sensitive. The same peptide, for example, can produce hemolysis ratios that vary by up to four‑fold when the assay is run on mouse versus rabbit erythrocytes, and simply substituting the positive‑control detergent can alter the value by a further approximately three‑fold [24]. Our dataset comprises HC50 values measured with human, mouse, horse and rabbit erythrocytes, yet key protocol details are often left unspecified; consequently, substantial inter‑experiment variability is inevitable. To limit overfitting and ensure generalization across these heterogeneous conditions, a robust loss function is fundamentally essential. We evaluated Mean Square Loss (MSE), Mean Absolute Loss (MAE), Huber Loss and Log-Cosh Loss (Eq. 8). As expected, MSE performed worst because of its pronounced sensitivity to outliers. Huber Loss comes with an additional hyperparameter to tune to control the switch from quadratic and linear loss. We experimented with $\delta=0.5$ and $\delta=1.0$, with both giving better results than MAE, but still slightly trailed Log-Cosh. Log-Cosh seamlessly integrates quadratic and linear penalties into a single smooth, twice-differentiable function, eliminating the need to hand-tune a $\delta$ hyperparameter while dampening the influence of large residuals.
@@
TABLE IV: Hyperparameter search space for AmpLyze.
Hyperparameter Candidates Best
@@
$d_{\mathrm{model}}$ 60, 128, 200, 300 128
IV Conclusion
@@
In this work, we present AmpLyze, a sequence‑based deep learning model to predict peptide hemolytic concentration (HC50) quantitatively and interpretably. By leveraging pre-trained protein language model embeddings and a fusion‑embedding strategy, AmpLyze achieves superior performance over prior regression approaches: Pearson’s $r$ improved from 0.739 to 0.756, $R^{2}$ rose from 0.543 to 0.570, MSE fell from 1.074 to 0.987, and MAE declined from 0.753 to 0.703. These gains demonstrate that deep neural architectures can more effectively capture the nonlinear, high‑dimensional relationships between sequence motifs and hemolytic activity. We further show that robust loss functions are crucial for handling the inherent heterogeneity of hemolysis assays. Experimental HC50 values can vary by orders of magnitude across different erythrocyte sources, buffer compositions, and protocols. We found that losses such as Log‑Cosh enables AmpLyze to generalize across noisy, multi‑source datasets. Importantly, AmpLyze offers residue‑level interpretability via Expected Gradients with a multi‑baseline scheme anchored by non‑hemolytic references. This approach successfully recapitulates literature‑reported mutational effects, shedding light on rational peptide design.
@@
Looking ahead, the next step is to integrate AmpLyze with in-silico MIC predictors to create a unified HC50–MIC framework. Both hemolytic and antimicrobial activities depend on peptide–membrane interactions: increased cationicity and amphipathicity promote binding to negatively charged bacterial membranes but also enhance insertion into zwitterionic erythrocyte bilayers. Future work will focus on training a multi‑task model to capture subtle sequence features that maximize bacterial pore formation while minimizing red‑cell lysis. Such a joint optimization strategy promises to accelerate the design of peptides with balanced safety and efficacy.